# Subsystem localization

### v2

@@ Abstract
We consider a ladder system where one leg, referred to as the “bath”, is governed by an Aubry-André (AA) type Hamiltonian, while the other leg, termed the “subsystem”, follows a standard tight-binding Hamiltonian. We investigate the localization properties in the subsystem induced by its coupling to the bath. For the coupling strength larger than a critical value ($t^{\prime}>t^{\prime}_{c}$), the analysis of the static properties show that there are three distinct phases as the AA potential strength $V$ is varied: a fully delocalized phase at low $V$, a localized phase at intermediate $V$, and a weakly delocalized (fractal) phase at large $V$. An analysis of the wavepacket dynamics shows that the delocalized phase exhibits a ballistic behavior, whereas the weakly delocalized phase is subdiffusive. Interestingly, we also find a superdiffusive narrow crossover regime along the line separating the delocalized and localized phases. When $t^{\prime}<t^{\prime}_{c}$, the intermediate localized phase disappears, and we find a delocalized (ballistic) phase at low $V$ and a weakly delocalized (subdiffusive) phase at large $V$. Between those two phases, there is also a crossover regime where the system can be super- or subdiffusive. Finally, in some limiting scenario, we also establish a mapping between our ladder system and a well-studied one-dimensional generalized Aubry-André (GAA) model.
@@ I Introduction
Anderson localization is a key quantum effect where uncorrelated disorder localizes wavefunctions and blocks transport of non-interacting electrons [1, 2, 3]. It creates a mobility edge in three dimensions that separates localized and extended states. A metal-insulator transition occurs when the Fermi energy crosses this edge. In one dimension (1D), all states become localized with any amount of disorder. There is no mobility edge and no transition. Introducing interactions in such disordered Anderson localized systems can stabilize a many-body localized (MBL) phase, where systems fail to thermalize and retain memory of initial conditions [4, 5, 6, 7, 8, 9, 10, 11]. This behavior contradicts the Eigenstate Thermalization Hypothesis (ETH), which posits that isolated non-integrable systems should relax to thermal equilibrium [12, 13, 14, 15]. However, recent studies highlight the avalanche instability within the MBL phase. This instability suggests that rare regions of the system can thermalize and spread thermalization through the system, potentially destabilizing the MBL phase [16, 17, 18].
@@
On the other hand, the presence of correlated disorder can significantly change the Anderson localization scenario in 1D. One can find the localization-delocalization transition and even the existence of the mobility edge, which is also feasible in such systems. One of the prime examples of such systems is the Aubry-André (AA) Hamiltonian, where the uncorrelated ”true” disorder is replaced by a quasi-periodic potential  [19]. One can induce a delocalization–localization transition by tuning the strength of the quasi-periodic potential, a phenomenon that holds significance in various physical contexts  [20, 21, 22, 23, 24]. The AA Hamiltonian also has the unique property of self-duality at the critical point, with identical real space and momentum space representations, thus it contains an energy-independent localization transition for the whole spectrum with no mobility edges. It has been shown that by adding different types of perturbations to the AA model, the fine-tuned AA duality conditions can be broken, and an energy-dependent self-duality relation can be established, automatically implying the existence of the mobility edges [25, 26, 27, 28, 29, 30, 31, 32]. With the remarkable progress in cold-atom and ion-trap experiments over the past decade, many of these models have been experimentally realized [33, 34, 35, 36], providing strong motivation to study quasi-periodic systems beyond purely theoretical interest. In this context, the emergence of a non-ergodic metallic phase has been proposed in models with single-particle mobility edges when interactions are introduced. [37, 38, 39, 40, 41]. One of the questions in the field of localization has persisted for some time now, i.e., what happens if one couples localized and extended systems together? In the case of a many-body interacting system, if the degrees of freedom of the extended system are large enough compared to the localized system, the extended system is expected to thermalize the localized one. On the other hand, if the delocalized degrees of freedom are comparable to the localized ones, the outcome is not yet clear. Many efforts have been made to address this question in the last several years [42, 43, 44, 45, 46]. Numerical challenges to tackle the exponentially growing Hilbert space dimension with system size and limitations of analytical tools to solve interacting systems have made this problem extremely challenging. This manuscript has tried to address a similar question in the non-interacting context.
@@
We consider a ladder system where one leg, referred to as the “bath” $B$ is a 1D non-interacting disordered chain, while the other leg, termed the “subsystem” $A$, is described by a clean delocalized tight-binding Hamiltonian (see the schematic of our system in Fig. 1). If the disorder in bath $B$ is uncorrelated, according to the rule of Anderson localization (i.e., any tiny disorder is sufficient to localize all states in 1D and 2D), both the subsystem $A$ and Bath $B$ of the total system are expected to be Anderson localized. However, if the disorder is correlated, the outcome is not at all apparent, given that the so-called rule of Anderson localization does not apply there. We investigate such a scenario by introducing a quasi-periodic potential in bath $B$. To be more precise, our $B$ part of the system is described by the AA Hamiltonian (which undergoes a delocalization-localization transition with increasing the strength of the quasi-periodic potential). We find that, depending on the strength of the quasi-periodic potential $V$ and the coupling strength $t^{\prime}_{AB}$ between subsystem $A$ and bath $B$, it is possible to localize the subsystem. One might expect such a situation to arise (if at all) in the large $V$ limit, when the bath $B$ is strongly localized. Interestingly, it turns out that is not the case. In contrast, this situation arises when $t^{\prime}_{AB}$ is larger than a critical strength $t^{\prime}_{c}$, and that too for an intermediate range of $V$. In the $V\gg t^{\prime}_{AB}$ limit, the subsystem A remains (weakly) delocalized and displays subdiffusive transport.
@@
Figure 1: Schematic diagram of the model: the subsystem ($A$) is described by a standard TB Hamiltonian whereas the bath ($B$) is governed by an AA type potential.
@@ II.1 Model system
In this paper, we investigate a ladder system whose one leg is called the “subsystem” and the other leg is called the “bath”. These are, respectively, indicated by “$A$” and “$B$” in the schematic diagram (Fig. 1). The subsystem $A$ is described by a standard tight-binding (TB) Hamiltonian ($H_{A}$ in Eq. 1) while the bath is governed by an Aubry-André (AA) type Hamiltonian ($H_{B}$ in Eq. 1). Without the coupling ($t^{\prime}=0$) between the subsystem $A$ and bath $B$, all the states in $A$ are delocalized. The problem we investigate here is whether and when states become localized in $A$ as we establish a coupling ($t^{\prime}\neq 0$) between $A$ and $B$.
Figure 2: Schematic phase diagram: P1 - delocalized (ballistic) phase, P2 - localized phase, P3 - weakly delocalized (fractal and subdiffusive) phase, and P4 - a crossover regime (above $t^{\prime}_{c}$ it is superdiffusive, and below $t^{\prime}_{c}$ it is super- or subdiffusive).
@@ II.2 Phase Diagram
The summary of our main results is presented in the schematic diagram of Fig. 2. The study of the static properties, as discussed in Sec. III, reveals that above a certain threshold value of the coupling strength ($t^{\prime}>t^{\prime}_{c}$), we have three distinct phases - a fully delocalized phase (for small values of $V$), a fully localized phase (for intermediate values of $V$) and a weakly delocalized (fractal) phase (for large values of $V$). We see a narrow crossover region with fractal character. The study of wavepacket dynamics, as discussed in Sec. IV, reveals that the delocalized phase is ballistic and the weakly delocalized phase is subdiffusive in nature. Interestingly, the crossover region is found to be superdiffusive in nature.
In the regime where $t^{\prime}<t^{\prime}_{c}$, the study of static properties shows the absence of the localized phase. Here, one finds two phase - a delocalized phase for small values of $V$ and a weakly delocalized (fractal) phase for large value of $V$. The study of wavepacket dynamics reveals that there is some intermediate crossover regime which is a fractal phase but can be super- or subdiffusive in nature.
It may be worth noting here that the phase diagram presented here strongly depends on the nature of the potential in the bath $B$. Instead of correlated AA type potential, if we take a random potential (for example, replace the uniform phase factor $\phi$ in Eq. 1 by a site-dependent uncorrelated phase factor $\phi_{i}$), the subsystem A will always be in a localized phase, in accordance with the Anderson localization in a disordered one-dimensional system. The relevant result is presented in Fig. 13.
@@ III Static properties
In this section, we study the localization properties of the subsystem $A$, as a function of the coupling strength $t^{\prime}$ and the AA potential strength $V$. For this purpose, we calculate the Inverse Participation Ratio for the subsystem (denoted by $IPR_{A}$) and the average Participation Ratio for the subsystem (denoted by $\langle PR_{A}\rangle$). To evaluate $IPR_{A}$, we first project an eigenket of the full Hamiltonian (Eq. 1) onto the subsystem $A$. We then calculate $IPR_{A}$ from the normalized projected state. If |$\ket{\xi_{n}}$ is the $n$th eigenket of the full Hamiltonian, and $\hat{P}_{A}$ is the projector for the subsystem $A$, the corresponding normalized projected state for the subsystem is:
$|\psi_{n}\rangle=\frac{\hat{P}_{A}|\xi_{n}\rangle}{||\hat{P}_{A}|\xi_{n}\rangle||}=\frac{\sum_{j=1}^{N}\langle j|\xi_{n}\rangle|j\rangle}{||\sum_{j=1}^{N}\langle j|\xi_{n}\rangle|j\rangle||}.$ (2)
It may be noted here that we choose the standard site basis {|$\{\ket{j}\}$ for our calculations, where |$\ket{j}$ represents a ket corresponding to a single particle at the $j$th site (and no particles at any other site). We define the subsystem IPR in the following way:
$\mathrm{IPR}_{A}^{(n)}=\sum_{j\in A}|\braket{j}{\xi_{n}}|^{4},$ (3)
where $n$ is the eigenket index.
It is worth noting here that all the eigenkets of the full Hamiltonian do not have the same degree of significance when studied from the subsystem $A$. To study the localization on the subsystem $A$, we only analyze those eigenkets for which the subsystem probability $p_{A}$ is larger than certain value ($\epsilon$), where
$p_{A}=\sum_{j\in A}|\braket{j}{\xi}|^{2},$ (4)
for an eigenket |$\ket{\xi}$ of the full Hamiltonian. An eigenket is called an $\epsilon$-significant if $p_{A}\geq\epsilon$. For ϵ=italic-$\epsilon=$ 0.3, 0.5 and 0.7, we calculate and analyze $IPR_{A}$ and $\langle PR_{A}\rangle$ (defined later) of the $\epsilon$-significant eigenkets.
@@
Figure 5: Variation of $V_{1}$ and $V_{2}$ as function of $t^{\prime}$.
To see how many eignkets are $\epsilon$-significant for the subsystem $A$, we define a parameter $\nu$ in the following way. Let $\mathcal{S}_{\epsilon}=\{\ket{\xi}~{}:p_{A}\geq\epsilon\}$ be the set of $\epsilon$-significant eigenkets, i.e., the set of eigenkets for which the subsystem probability $p_{A}\geq\epsilon$. Since there are $2N$ number of eigenkets, we define:
@@
Figure 7: Log-log plots: $\langle PR_{A}\rangle$ vs. N for $V=1,5,10,17$ ($\epsilon=0.5$ and $t^{\prime}=5$). Inset: Scaling exponent ($\eta$) vs. V.

### v3

@@ Abstract
We consider a ladder system where one leg, referred to as the “bath”, is governed by an Aubry-André (AA) type Hamiltonian, while the other leg, termed the “subsystem”, follows a standard tight-binding Hamiltonian. We investigate the localization properties in the subsystem induced by its coupling to the bath. For the coupling strength larger than a critical value ($t^{\prime}>t^{\prime}_{c}$), the analysis of the static properties shows that there are three distinct phases as the AA potential strength $V$ is varied: a fully delocalized phase at low $V$, a localized phase at intermediate $V$, and a weakly delocalized (fractal) phase at large $V$. The fractal phase also appears in a narrow region along the boundary between the delocalized and localized phases. An analysis of the projected wavepacket dynamics in the subsystem shows that the delocalized phase exhibits a ballistic behavior, whereas the weakly delocalized phase is subdiffusive. Interestingly, the narrow fractal phase shows a super- to subdiffusive behavior as we go from the delocalized to localized phase. When $t^{\prime}<t^{\prime}_{c}$, the intermediate localized phase disappears, and we find a delocalized (ballistic) phase at low $V$ and a weakly delocalized (subdiffusive) phase at large $V$. Between those two phases, there is also an anomalous crossover regime where the system can be super- or subdiffusive. Beyond the ballistic phase observed at low $V$, we also identify a superdiffusive regime emerging in the limit $t^{\prime}/V\ll 1$, which continuously approaches the ballistic behavior as $t^{\prime}\to 0$. Finally, in some limiting scenario, we also establish a mapping between our ladder system and a well-studied one-dimensional generalized Aubry-André (GAA) model.
@@ I Introduction
Anderson localization is a key quantum effect where uncorrelated site disorder localizes wavefunctions and blocks transport of non-interacting electrons [1, 2, 3]. It creates in a three dimensional system a mobility edge that separates localized and extended states. A metal-insulator transition occurs when the Fermi energy crosses this edge. In one dimension (1D), all states become localized with any amount of disorder. There is no mobility edge and no transition. Introducing interactions in such disordered Anderson localized systems can stabilize a many-body localized (MBL) phase, where systems fail to thermalize and retain memory of initial conditions [4, 5, 6, 7, 8, 9, 10, 11]. This behavior contradicts the Eigenstate Thermalization Hypothesis (ETH), which posits that isolated non-integrable systems should relax to thermal equilibrium [12, 13, 14, 15]. However, recent studies highlight the avalanche instability within the MBL phase. This instability suggests that rare regions of the system can thermalize and the growth of entanglement between different such regions potentially destabilizing the MBL phase [16, 17, 18].
@@
On the other hand, the presence of correlated disorder can significantly change the Anderson localization scenario in 1D. Such systems may have a localization-delocalization transition and even may have a mobility edge. One of the prime examples of such systems is the Aubry-André (AA) Hamiltonian, where the uncorrelated “true” disorder is replaced by a quasi-periodic potential  [19]. One can induce a delocalization–localization transition by tuning the strength of the quasi-periodic potential, a phenomenon that holds significance in various physical contexts  [20, 21, 22, 23, 24]. The AA Hamiltonian also has the unique property of self-duality at the critical point, with identical real space and momentum space representations, thus it contains an energy-independent localization transition for the whole spectrum with no mobility edges. It has been shown that by adding different types of perturbations to the AA model, the fine-tuned AA duality conditions can be broken, and an energy-dependent self-duality relation can be established, automatically implying the existence of the mobility edges [25, 26, 27, 28, 29, 30, 31, 32]. With the remarkable progress in cold-atom and ion-trap experiments over the past decade, many of these models have been experimentally realized [33, 34, 35, 36], providing strong motivation to study quasi-periodic systems beyond purely theoretical interest. In this context, the emergence of a non-ergodic metallic phase has been proposed in models with single-particle mobility edges when interactions are introduced [37, 38, 39, 40, 41].
A question that has persisted for some time now is what happens if one couples localized and extended systems together? In the case of a many-body interacting system, if the degrees of freedom of the extended system are large enough compared to the localized system, the extended system is expected to thermalize the localized one. On the other hand, if the delocalized degrees of freedom are comparable to the localized ones, the outcome is not yet clear. Many efforts have been made to address this question in the last several years [42, 43, 44, 45, 46]. Numerical challenges to tackle the exponentially growing Hilbert space dimension with system size and limitations of analytical tools to solve interacting systems have made this problem extremely challenging. In this work, we address a similar question in the non-interacting context.
@@
We consider a ladder system where one leg, referred to as the “bath” is a 1D non-interacting disordered chain, while the other leg, termed the “subsystem”, is described by a clean delocalized tight-binding (TB) Hamiltonian (see the schematic of our system in Fig. 1). If the disorder in bath is uncorrelated, according to the rule of Anderson localization (i.e., any tiny disorder is sufficient to localize all states in 1D and 2D), both the subsystem and bath of the total system are expected to be Anderson localized. However, if the disorder is correlated, the outcome is not at all apparent, given that the so-called rule of Anderson localization does not apply there. We investigate such a scenario by introducing a quasi-periodic potential (AA type) in the bath. We find that, depending on the strength of the quasi-periodic potential $V$ and the coupling strength $t^{\prime}$ between subsystem and bath, it is possible to localize the subsystem. One might expect such a situation to arise (if at all) in the large $V$ limit, when the bath is strongly localized. Interestingly, it turns out that is not the case. In contrast, this situation arises when $t^{\prime}$ is larger than a critical strength $t^{\prime}_{c}$, and that too for an intermediate range of $V$. In the limit of large $V$, the subsystem remains weakly delocalized and exhibits subdiffusive transport. Finally, in some limiting scenario, we also establish a mapping between our system and a well-studied one-dimensional generalized Aubry-André (GAA) model, which also exhibits similar ballistic, superdiffusive, and subdiffusive phases, as reported recently [47].
Manuscript is organized as follows. In Sec. II, we discuss our model and outline the main outcome in the form of a phase diagram. In Sec . III, we investigate the static properties of the subsystem for the significant states. Sec. IV is dedicated to studying the projected wave-packet dynamics in the subsystem. In Sec. V, we discuss the analytical understanding of our results. Finally, in Sec.VI, we briefly summarize our main conclusions.
@@
Figure 1: Schematic diagram of the model: the subsystem (AA) is described by a standard TB Hamiltonian whereas the bath ($B$) is governed by an AA type potential.
@@ II.1 Model system
In this paper, we investigate a ladder system whose one leg is called the “subsystem” and the other leg is called the “bath”. These are, respectively, indicated by “AA” and “BB” in the schematic diagram (Fig. 1). The subsystem $A$ is described by a standard tight-binding (TB) Hamiltonian ($H_{A}$ in Eq. 1) while the bath is governed by an Aubry-André (AA) type Hamiltonian ($H_{B}$ in Eq. 1). Without the coupling ($t^{\prime}=0$) between the subsystem $A$ and bath $B$, all the states in $A$ are delocalized. The problem we investigate here is whether and when states become localized in $A$ as we establish a coupling ($t^{\prime}\neq 0$) between $A$ and $B$.
Figure 2: Schematic phase diagram in rescaled parameter space of $\tilde{t^{\prime}}=t^{\prime}/t^{\prime}_{c}$ and $\tilde{V}=V/t^{\prime}_{c}$. P1 - delocalized (ballistic) phase, P2 - localized phase, P3 - weakly delocalized (fractal and subdiffusive) phase, P4 - an anomalous crossover regime (fractal and super- or sub-diffusive), and P5 - weakly delocalized (fractal and superdiffusive) phase.
@@ II.2 Phase Diagram
The summary of our main results is presented in the schematic diagram of Fig. 2. It is constructed based on two scaling parameters: $\eta$ - finite-size scaling exponent of the average subsystem PR (Participation Ratio) and $\gamma$ - dynamical exponent in the long-time behavior of the MSD (Mean-Squared Displacement). The analysis of the static properties, as discussed in Sec. III, reveals that above a certain threshold value of the coupling strength ($t^{\prime}>t^{\prime}_{c}\approx 4.4$), we have three distinct phases - a fully delocalized phase (for small values of $V$), a fully localized phase (for intermediate values of $V$) and a weakly delocalized (fractal) phase (for large values of $V$). We also see a narrow crossover region with fractal character along the line separating the delocalized and localized phases. The study of projected wavepacket dynamics in subsystem, as discussed in Sec. IV, reveals that the delocalized phase is ballistic (referred to as P1 phase) and the weakly delocalized phase is subdiffusive (referred to as P3 phase) in nature. Interestingly, the narrow crossover region (referred to as P4 phase) is found to exhibit both super- and sub-diffusive behavior as we go across the region.
In the regime where $t^{\prime}<t^{\prime}_{c}$, the study of static properties shows the absence of the localized phase. Here, one finds two phases - a delocalized phase for small values of $V$ and a weakly delocalized (fractal) phase for large values of $V$, which also shows ballistic and sub-diffusive transport, and hence, identified as once again P1 and P3, respectively. The study of wavepacket dynamics reveals that P4 phase also exists for $t^{\prime}<t^{\prime}_{c}$ in some intermediate regime of $V$, which is a fractal phase but can be super- or subdiffusive in nature. Moreover, we also identify an additional superdiffusive regime, referred to as the P5 phase. Such a phase exists in the regime $t^{\prime}/V\ll 1$, and this phase becomes ballistic in nature in the strict limit $t^{\prime}\to 0$.
The schematic phase diagram (Fig. 2) summarizes all the main results of this work. For the phase diagram, we rescale the axes as $\tilde{t^{\prime}}=t^{\prime}/t^{\prime}_{c}$ and $\tilde{V}=V/t^{\prime}_{c}$, and restrict our regime of interest to $\tilde{t^{\prime}}\in[0,2]$ and $\tilde{V}\in[0,10]$. Since it is possible that, for a certain regime, a finite fraction of the significant states are localized and the rest are delocalized (weakly delocalized), it is important to have clarification on the meaning of the exponent $\eta$ in such a case and related phases in our energy-independent phase diagram. As we discuss later (see Sec. III and Appendix C), in the thermodynamic limits, all the significant states are delocalized (localized) in the P1 (P2) phase. In the fractal phase (P3, P4, and P5 phases together), a finite fraction of significant states are delocalized (weakly delocalized) in general, and the rest are localized. The scaling exponent $\eta$ is effectively determined by the average behavior of the delocalized (weakly delocalized) states. In this case, where we have both types of states, the phase diagram is determined by the nature of non-localized states since the dynamical nature of the subsystem (characterized by the exponent $\gamma$) is mainly dependent on these states.
It may be noted that the phase diagram presented here strongly depends on the nature of the potential in the bath $B$. Instead of a correlated AA-type potential, if we take a random potential (for example, replace the uniform phase factor $\phi$ in Eq. 1 by a site-dependent uncorrelated phase factor $\phi_{i}$), the subsystem $A$ will always be in a localized phase, in accordance with the Anderson localization in a disordered one-dimensional system. The relevant result is presented in Fig. 16 of Appendix A. A similar model with true disorder was recently examined in Ref. [48], though there the authors focused on the entire system, unlike our analysis of the subsystem.
@@ III Static properties
In this section, we study the localization properties of the subsystem $A$, as a function of the coupling strength $t^{\prime}$ and the AA potential strength $V$. For this purpose, we calculate the Inverse Participation Ratio for the subsystem (denoted by $IPR_{A}$) and the average Participation Ratio for the subsystem (denoted by $\langle PR_{A}\rangle$). Before we discuss about these measures and corresponding results, we first note that all the eigenkets of the full Hamiltonian do not have the same degree of significance when studied from the subsystem $A$. We, therefore, begin our discussion with an analysis on the significant states.
@@
Figure 5: Variation of $V_{1}$ and $V_{2}$ as function of $t^{\prime}$. The extrapolation of data shows that $V_{1}\approx V_{2}$ at $t^{\prime}=t^{\prime}_{c}\approx 4.4$ (represented by a star symbol).
III.1 The $\epsilon$-significant states
All the eigenkets of the full Hamiltonian do not have the same degree of significance when studied from the subsystem $A$. To study the localization on the subsystem $A$, we only analyze those eigenkets for which the subsystem probability $p_{A}$ is larger than certain value ($\epsilon$), where
$p_{A}=\sum_{j\in A}|\braket{j|\xi}|^{2},$ (2)
[…]
@@ III.1 The $\epsilon$-significant states
III.2 $IPR_{A}$: Inverse Participation Ratio for subsystem
[…]
$|\psi_{n}\rangle=\frac{\hat{P}_{A}|\xi_{n}\rangle}{\lVert\hat{P}_{A}|\xi_{n}\rangle\rVert}=\frac{\sum_{j\in A}\langle j|\xi_{n}\rangle|j\rangle}{\lVert\sum_{j\in A}\langle j|\xi_{n}\rangle|j\rangle\rVert}.$ (4)
We define the subsystem $IPR_{A}$ in the following way:
$\mathrm{IPR}_{A}^{(n)}=\sum_{j\in A}|\braket{j|\psi_{n}}|^{4},$ (5)
where $n$ is the eigenket index.
@@ III.2 $IPR_{A}$: Inverse Participation Ratio for subsystem
Figure 7: Log-log plots: $\langle PR_{A}\rangle$ vs. N for $V=1,5,10,17$ ($\epsilon=0.5$ and $t^{\prime}=5$). The black dashed lines show the fitting to the functional form $\langle PR_{A}\rangle\sim N^{\eta}$.
Inset: Scaling exponent $\eta$ vs. V.